# Where Quantum Complexity Helps Classical Complexity: A Survey

### v2

@@ Abstract.
Scientists have demonstrated that quantum computing has presented novel approaches to address computational challenges, each varying in complexity. Adapting problem-solving strategies is crucial to harness the full potential of quantum computing. Nonetheless, there are defined boundaries to the capabilities of quantum computing. This paper concentrates on aggregating prior research efforts dedicated to solving intricate classical computational problems through quantum computing. The objective is to systematically compile an exhaustive inventory of these solutions and categorize a collection of demanding problems that await further exploration.
@@ 1. Introduction
On the other hand, Quantum Computing utilizes unique units called qubits, which can be in states of 0, 1, or a combination of both due to quantum properties. It harnesses quantum phenomena, such as superposition and entanglement, to conduct multiple computations simultaneously, potentially leading to exponential acceleration for specific problems. Quantum computing is rooted in the principles of quantum mechanics, allowing qubits to exist in multiple states and interact through entanglement. Information is stored in qubits, but their susceptibility to quantum decoherence poses a challenge. Quantum computing has the potential to deliver exponential speedup for particular problems through parallelism and quantum interference, although this benefit is not universal. Presently, quantum computing is in its nascent stage, predominantly focusing on specialized tasks. Practical quantum computers are not as widespread as classical computers. Anticipations arise regarding the capacity of quantum computers to effectively tackle complex problems involving a multitude of variables and potential outcomes at a much swifter pace compared to classical computers.
@@
Numerous prior studies have focused on addressing challenging problems in classical computing. We have compiled a collection of these articles to identify unresolved and significant problems worthy of attention.
The overview of these papers is provided in Section 2. In Section 3, we individually outline and provide concise descriptions of each of these works. The classification of open problem findings from these preceding research endeavors is the focal point of Section 5.
2. Summary
@@
In recent years, quantum computing has emerged as a promising avenue for addressing complex computational challenges. A series of notable studies have explored the application of quantum algorithms to a variety of combinatorial optimization problems, ranging from graph theory (de2022quantum, ; mariella2023quantum, ; siraichi2019qubit, ; li2020qubit, ; bravyi2022hybrid, ; bochniak2023quantum, ; salehi2022unconstrained, ; Ge_2020, ; cocchi2021graph, ; pichler2018quantum, ; kim2022rydberg, ; doern2007quantum, ; 8477865, ; pelofske2019solving, ; kerger2023asymptotically, ; pelofske2019solving, ; ramanujan2021approximate, ; kimmel2021query, ; witter2022query, ; dorn2009quantum, ; gabow1990data, ; mcleod2022benchmarking, ; vert2020revisiting, ; cameron2006quantum, ; bravyi2022hybrid, ; bravyi2020obstacles, ) to chemistry simulations (kassal2011simulating, ).
@@
Researchers are exploring how quantum computers can solve complex problems. They convert 3-SAT problems into a form that quantum computers can handle better. An improved method called ab-QAOA (yu2023solution, ) helps solve hard 3-SAT and Max-3-SAT problems using fewer steps. Quantum annealers, like D-Wave, are tested to solve SAT and MaxSAT problems. A bio-molecular (chang2021quantum, ) algorithm is proposed to solve the independent set problem efficiently. QAOA’s behavior on graphs is studied, revealing its performance patterns. A quantum algorithm is developed to speed up solving the Traveling Salesman Problem (TSP). The Grover algorithm is used to tackle the Hamiltonian Cycle Problem (HCP) (jiang2022quantum, ).
@@
Researchers have been exploring quantum computing techniques to solve challenging optimization problems, such as the Multidimensional Knapsack Problem. They combined Quantum Particle Swarm Optimization with local search methods, yielding competitive outcomes. Another approach used the Quantum Approximate Optimization Algorithm (QAOA) to enhance solutions for constrained optimization problems, outperforming classical methods. The integration of the k-nearest neighbor technique and quantum cuckoo search led to better results in resource allocation and knapsack problems. A Diversity-Preserving Quantum Particle Swarm Optimization strategy (lai2020diversity, ) managed populations efficiently for improved solutions.
@@
The Quantum Alternating Operator Ansatz (cook2020quantum, ) offered approximate solutions for combinatorial problems, while the Quantum Approximate Optimization Algorithm tackled minimum vertex cover problems (pelofske2019solving, ; zhang2022applying, ; wang2023quantum, ; marsh2019quantum, ; wang2022classically, ). The Grover search algorithm achieved quantum speedup for the vertex cover problem. A Quantum-Walk assisted algorithm targeted bounded NP optimization problems. Notably, quantum algorithms were designed for matching problems, proving their effectiveness in various scenarios (jeffery2022multidimensional, ; bonnetain2020improved, ; mahasinghe2019solving, ; kempe2003quantum, ; childs2009universal, ; lovett2010universal, ; apers2019unified, ; marsh2019quantum, ; belovs2012learning, ; jeffery2022multidimensional, ).
Researchers have been exploring the capabilities of quantum computing in solving complex optimization problems. McLeod and Sasdelli (mcleod2022benchmarking, ) benchmarked D-Wave’s quantum annealers, Advantage, and 2000Q, using the maximum cardinality matching problem. They found that Advantage outperforms 2000Q for larger problems and analyzed the scaling of diabatic transition probability. Fowler addressed the minimum weight perfect matching problem on a modular qubit array, proving efficient solutions even in the presence of errors.
@@
In the realm of quantum walks, Jeffery and Zur (jeffery2022multidimensional, ) improved upper bounds on the time complexity of $k$-distinctness problems, offering insights into the quantum speedup achieved in these scenarios.
@@
Harris and Kashefi (harris2023scalable, ) introduced an error mitigation protocol. The protocol is for running BQP-type computations on a quantum computer s.t. the noises are time-dependent. The authors hired verification techniques by Leichtle et al. (leichtle2021verifying, ). This allowed their results to interleave standard and test computation rounds. One should notice that the above technique provides a local correctness guarantee. In particular, it bounds the probability of returning a correct classical output exponentially. The authors further introduced a post-selection technique they named basketing. It could address time-dependent noise behaviors and also allow overhead reduction.
@@
2.1. Statistical Analysis
Table 1 presents a comparison of time complexities for various problems in the context of Quantum versus Classical computing. The table consists of two columns: ”Classic C.” and ”Quantum C.”, each providing information based on the latest level of our knowledge in the field.
There is an additional column labeled ”Improvement.” This column aims to clarify whether a Quantum solution can be considered an improvement for a given problem. If a quantum computer solves a problem in time $T$ and a classical computer requires a polynomial function of $T$ (such as $T^{2}$), it is termed a polynomial speedup. When the time complexity of the best classical algorithm, denoted as $C(n)$, and the time complexity of the quantum algorithm, denoted as $Q(n)$, satisfies $C=2^{\Omega(Q^{\lambda})}$ for some positive constant $\lambda$, the speedup is referred to as superpolynomial. Exponential speedup occurs when a quantum computer requires time $T$, while a classical computer needs an exponentially growing function of $T$ (for example, $2^{T}$). These categories represent the promising and desired speedups in quantum computers.
By Problem Number(s) in the last column, we refer to the corresponding open problem(s) mentioned in the paper. In this table, $|V|=n$ is the number of vertices and $|E|$ is the number of edges in a graph.
@@ 2.1. Statistical Analysis
Problem Classic C. Quantum C. Improvement Problem Number(s) k-distinctness $\Theta(n\log n)$ $\tilde{\mathcal{O}}(n^{\frac{3}{4}-\frac{1}{4}\frac{1}{2^{k}-1}})$ Polynomial - (Element Distinctness) (In decision tree model (10.1145/800061.808735, ))  (jeffery2022multidimensional, ) k-SAT ($k\geq 3$) $\mathcal{O}(2^{1-\frac{\Theta(1)}{k}})$ (10.1145/3313276.3316359, ) $\mathcal{O}((2(1-1/k))^{n})$ (schoning1999probabilistic, ) - 1, 2, 15, 16, 17, 66 32, 33, 34, 35, 36 Maximum Independent Set $\mathcal{O}(1.1996^{n}\cdot n^{\mathcal{O}(1)})$ (XIAO2017126, ) $\mathcal{O}(1.1488^{n})$ (etde_21241168, ) Polynomial 67, 98, 104, 98 Hamiltonian Cycle $\mathcal{O}^{*}(1.657^{n})$ (bjorklund2014determinant, ) $\mathcal{O}(n^{2n/(n+1)})$ (dorn2007quantum, ) Exponential 18, 10, 11, 14 Eulerian Tour $\mathcal{O}(|E|)$ (fleischner1990eulerian, ) $\mathcal{O}(\sqrt{n})$ (dorn2007quantum, ) Polynomial - Travelling Salesman Problem $\mathcal{O}(n^{2}\cdot 2^{n})$ (doi:10.1137/0110015, ) $\mathcal{O}({1.728}^{n}\cdot\text{poly}(n))$ (doi:10.1137/1.9781611975482.107, ) - 14 Maximum Matching $\mathcal{O}(|E|\sqrt{|V|})$ (micali1980v, ) $n^{3/2}\log^{2}n$ (beigi2022time, ) Polynomial 68, 86, 69, 71, 103 Maximum Cut $\text{exp}(n)$ (garey1997computers, ) $\sqrt{2^{n}/r}$ (chang2023quantum, ) - 8, 6 s.t. $r$ is num of max-cuts Graph Isomorphism $2^{\mathcal{O}(\log n)^{c}}$ (grohe2020graph, ) NA (alagic2007quantum, ) NA 96, 79, 31, 58 Maximum Clique $2^{n/4}$ (robson2001finding, ) $\mathcal{O}(\log_{2}n\cdot\sqrt{2^{n}})$ (haverly2021comparison, ) - 88, 89 String Edit Distance $\mathcal{O}(s^{2}+n)$ (landau1998incremental, ) $\mathcal{O}(n^{2}/\log n)$ (equi2021quantum, ) Poly-logarithmic 74, 91, 92 s.t. $s$ is a max. edit distance Longest Common Subsequence $\mathcal{O}(n\log^{2}n)$ (bhowmick2019approach, ) $\tilde{\mathcal{O}}(\sqrt{n})$ (le2023quantum, ) Polynomial 42 Protein Folding $\omega(n^{c}),\ \forall c>0$ $\mathcal{O}(2^{n/2})$ (wong2022fast, ) - 27, 101, 102 Mixed-Integer Programming $(\log n)^{{}^{\mathcal{O}(n)}}\cdot\text{poly}(n)$ (reis2023subspace, ) $\exp(n)$ 13 Knapsack Problem $\mathcal{O}(2^{n/2})$ (doi:10.1137/0210033, ) $\mathcal{O}(2^{n/2})$ (fujimura2010quantum, ) - 39, 38, 5.12, 41 Subset-Sum $\tilde{\mathcal{O}}(2^{0.291n})$ (becker2011improved, ) $\tilde{\mathcal{O}}(2^{0.236n})$ (bonnetain2020improved, ) - 37, 44, 45, 46 Discrete Logarithm $\omega(n^{c}),\ \forall c>0$ $\mathcal{O}(\log|\mathcal{G}|)$ (hhan2023quantum, ) Superpolynomial - for a cyclic group $\mathcal{G}$ of prime order Integer Factorization $\exp(1.9(\log n)^{\frac{1}{3}}(\log\log n)^{\frac{2}{3}})$ (buhler1993factoring, ) ${\mathcal{O}}(b^{3})$ (shor1999polynomial, ) Superpolynomial - $n$ is the number to factor s.t. $b$ is num of bits of $n$ Crypto. Certified Random Bits - $\exp(n)$ (aaronson2023certified, ) - 50, 55

### v4

@@ Abstract.
Scientists have demonstrated that quantum computing has presented novel approaches to address computational challenges, each varying in complexity. Adapting problem-solving strategies is crucial to harness the full potential of quantum computing. Nonetheless, there are defined boundaries to the capabilities of quantum computing. This paper concentrates on aggregating prior research efforts dedicated to solving intricate classical computational problems through quantum computing. The objective is to systematically compile an exhaustive inventory of these solutions and categorize a collection of demanding open problems that await further exploration.
@@ 1. Introduction
Quantum Computing utilizes unique units called qubits, which can be in states of 0, 1, or a combination of both due to quantum properties. It harnesses quantum phenomena, such as superposition and entanglement, to conduct multiple computations simultaneously, potentially leading to exponential acceleration for specific problems. Quantum computing is rooted in the principles of quantum mechanics, allowing qubits to exist in multiple states and interact through entanglement. Information is stored in qubits, but their susceptibility to quantum decoherence poses a challenge. Quantum computing has the potential to deliver exponential speedup for particular problems through parallelism and quantum interference, although this benefit is not universal. Presently, quantum computing is in its nascent stage, predominantly focusing on specialized tasks. Practical quantum computers are not as widespread as classical computers. Anticipations arise regarding the capacity of quantum computers to effectively tackle complex problems involving a multitude of variables and potential outcomes at a much swifter pace compared to classical computers.
@@
In contrast to classical systems, where the influence of the environment on the system is minimal, quantum systems can rapidly and significantly become entangled with their environment, leading to disturbances. Complete isolation of a quantum system from its environment is nearly impossible, making interaction between the two unavoidable. The phenomenon of quantum decoherence poses a challenge in describing the dynamics of quantum systems. Essentially, decoherence characterizes the effects of entangling interactions on the future measurement statistics of the system (SCHLOSSHAUER20191, ).
Quantum decoherence greatly hinders quantum information processing, such as quantum computation. Decoherence can impact the superposition states of quantum systems, leading to errors. To address this, qubit systems have been engineered to minimize the influence of environmental interactions on state preparation and longevity. Various strategies have been developed to mitigate the adverse effects of decoherence, ranging from preventing or minimizing errors (e.g., encoding information in a decoherence-free subspace) to correcting undesired changes (referred to as error correction) (SCHLOSSHAUER20191, ).
The quantum error correction (QEC) process is vital for addressing the decoherence problem. However, classical error correction algorithms cannot be easily applied to quantum systems due to some obstacles. The no-cloning theorem, as described in (wootters2009no, ), states that no machine can copy any given quantum state (note that some but not all states can be cloned). Additionally, measuring a quantum system disturbs the state and might destroy superposition. This means reading out information without destroying it is more complex in quantum systems than in classical ones. Quantum Noise has several continuous parameters, similar to some classical noises, but its probabilistic nature makes it harder to describe than the deterministic evolution of classical noise (lidar2013quantum, ).
Over time, researchers developed many QEC codes to mitigate the detrimental effects of decoherence. One of the famous codes is Peter Schor’s 9-qubit code, first introduced in 1995 in (1995PhRvA..52.2493S, ), which encodes $1$ logical qubit in $9$ qubits, capable of correcting arbitrary errors in a single qubit. However, there exist other codes with a higher rate, such as those developed by Calderbank, Shor (PhysRevA.54.1098, ), and Steane (1996, ) (referred to as CSS code) that were inspired by the classical Hamming code [4,7](paper (hamming1950errordetecting, ). Furthermore, a more general class of code known as stabilizer codes was discovered by Gottesman (gottesman1997stabilizer, ), and by Calderbank, Rains, Shor, and Sloane (calderbank1997quantum, ).
The measurement of quantum systems differs fundamentally from that in classical physics. In quantum systems, observables are measured experimentally, not the vector states. Consequently, without knowing the exact vector state, we cannot predict the outcome of the measurement. The state’s evolution during measurement is different from the evolution between measurements, contributing to the peculiar nature of this process. When an experimenter attempts to apply a specific observable operator on the system, the wave function collapses, meaning the state unpredictably jumps to one of the eigenstates of the observable operator (Susskind:2014qoa, ).
To achieve scalable quantum computation despite noise-process variations, quantum fault-tolerant methods are essential. Although there are similarities in the basic concepts of fault tolerance in classical and quantum computation, such as creating gadgets that encode information and correct errors simultaneously, and protecting the system from noise, implementing fault-tolerant computation in quantum systems is significantly more challenging owing to the coherent implementation of superposition states and the management of noise that can alter both the bit and the phase of the state (lidar2013quantum, ).
The distinction between quantum processors and classical ones lies in their utilization of quantum information instead of classical information. Quantum information possesses intrinsic features such as randomness, uncertainty, and entanglement, which are absent in classical systems. Quantum randomness alludes to an inherent randomness that does not stem from a lack of knowledge, unlike the everyday randomness we encounter. Uncertainty in the quantum system arises from the fact that certain observables cannot be measured simultaneously without interfering with each other, meaning that one measurement inevitably affects subsequent measurements. Furthermore, entanglement enables complete knowledge of the entire quantum system while imparting zero information about the individual subsystems (preskill2023quantum, ).
[…]
Different concepts of complexity in quantum computing involve considering the number of qubits used, analogous to storage in classical complexity, and the Kolmogorov complexity (berthiaume2001quantum, ), which measures the information content of a string and its compressibility without losing information. Quantum computational complexity can be centred on the number of operations, particularly significant in holographic duality and black hole interiors, offering a unique perspective distinct from traditional quantum computing. As the comprehension of the relationship between geometry and information advances, other complexity notions like Kolmogorov complexity in holography may also become pertinent (chapman2022quantum, ).
Numerous prior studies have focused on addressing challenging problems in classical computing by quantum approaches. Several others aimed to study such results, as Dalzell et al. (dalzell2310quantum, ) did. They extensively discussed the areas of applications of quantum solutions. Further, they thoroughly provided details about quantum algorithmic primitives, namely the building blocks to achieve quantum speedups. For each algorithmic primitive, they sketched its basic idea, example use cases and important caveats. They further evaluate potential quantum speedups by comparing quantum solutions to state-of-the-art classical methods and complexity-theoretic limitations. ZHU et al. (zhu2022brief, ) provided a brief survey of quantum architecture search methods that address the challenges of designing quantum circuit architectures for practical tasks such as combinatorial optimization. These methods are inspired by techniques from classical neural network architecture search and machine learning, such as reinforcement learning. The methods may also apply to other fields, such as multi-objective optimization. Valerio et al. (massoli2022leap, ) pointed out, in their survey work, that quantum neural networks and algorithms are still far from proving decisive supremacy over classical ones, yet it is an ongoing quest. Quantum technologies may later exploit phenomena such as superposition and entanglement to obtain improvements whatsoever. The authors discussed the latest achievements and state-of-the-art approaches to Quantum Perceptrons and Quantum Neural Networks. It is further elaborated that the Quantum machine learning algorithms, such as Quantum Neural Networks, accept quantum states as input, requiring the translation of classical data into quantum states through state preparation.
1.1. Overview of this Paper
In short, we have compiled a collection of articles to identify unresolved and significant problems worthy of attention. For numerous well-known problems in computer science, we have demonstrated a clear contrast between the quantum solutions’ runtimes and the classical ones. We explored the potential of quantum resources in achieving exponential speedup in computational tasks and the challenges of error correction and scalability in quantum systems. Moreover, we revisited hybrid quantum-classical algorithms that leverage the speed and efficiency of quantum processors while tackling the constraints associated with quantum systems. This emphasizes the importance of interdisciplinary collaboration in pioneering innovative solutions for classical computational challenges. We then clustered the directions possible for future investigations into several categories.
More specifically, form Section 2 to Section 6, we demonstrate the following:
Several notable studies have delved into the application of quantum algorithms to a diverse range of combinatorial optimization problems, including areas like graph theory (de2022quantum, ; mariella2023quantum, ; siraichi2019qubit, ; li2020qubit, ; bravyi2022hybrid, ; bochniak2023quantum, ; salehi2022unconstrained, ; Ge_2020, ; cocchi2021graph, ; kim2022rydberg, ; pelofske2019solving, ; kerger2023asymptotically, ; pelofske2019solving, ; ramanujan2021approximate, ; kimmel2021query, ; witter2022query, ; mcleod2022benchmarking, ; vert2020revisiting, ; bravyi2022hybrid, ; bravyi2020obstacles, ) and even chemistry simulations (kassal2011simulating, ). In section 2, we summarized the contributions of the previous studies, highlighting their methodologies and findings.
In section 3, as a descriptive analysis, we illustrated how bringing several classical problems to the quantum side showed theoretical and/or practical speedups in the literature. Further, we provided illustrations about where several well-known quantum computation techniques, e.g. Quantum Walk (apers2019unified, ; chailloux2021lattice, ; laarhoven2016search, ; marsh2019quantum, ; belovs2012learning, ; bernstein2013quantum, ) and Quantum Search (unsal2023faster, ; zhang2020procedure, ; srinivasan2018efficient, ; gheorghica2021gr3, ; miyamoto2020quantum, ; bonnetain2020improved, ), are applied. We specifically referenced results that applied these techniques to solve well-known classical problems such as the Hamiltonian Cycle Problem, the 3-SAT Problem, and others. Furthermore, we conducted an inferential analysis to illuminate more nuanced details of the current results and explore potential future directions to address existing open problems.
In section 4, we reviewed the existing findings on models of quantum complexity growth. We also revisited current results in the literature that illustrated the relationship between quantum randomness and classical complexity, showing how quantum randomness influences problem-solving. Furthermore, we delved into the approximation techniques in the literature for quantum problems, demonstrating how quantum algorithms can offer approximate solutions for complex computational tasks. Additionally, we covered quantum query complexity, that is, the efficiency of quantum algorithms in terms of query operations. We also revisited the concept of quantum random walk, highlighting its fundamental role in quantum algorithms and solving classical computational problems.
In Section 5, we reviewed various well-known computational problems that can be efficiently solved using quantum computing techniques. These include problems such as the Satisfiability Problem, Maximum Independent Set, Graph Traversal Problems, and Maximum Matching on Graphs. We revisited the existing results that showed how quantum algorithms can improve computational efficiency by leveraging quantum principles like superposition and entanglement. This further opens up new possibilities for effectively addressing challenging computational tasks.
In Section 6, we provided several open problems and potential areas for further exploration in the field of quantum computing. We covered various critical topics, including Quantum Approximate Optimization Algorithm (QAOA), Quantum Annealing, Quantum Search, Quantum Walk, and Mean Estimator. We also revisited Quantum String Matching Algorithms, Quantum Merging Trees, Quantum Congested Clique, Verification and Certification, Machine Learning, Complexity, Classification, Security, and related subjects. By highlighting these open problems, we aimed to assist future research endeavours toward enhancing quantum computing capabilities and addressing challenging computational problems.
2. Papers in Summary
@@
Especially in recent years, quantum computing has emerged as a promising avenue for addressing complex computational challenges. A series of notable studies have explored the application of quantum algorithms to a variety of combinatorial optimization problems, ranging from graph theory (de2022quantum, ; mariella2023quantum, ; siraichi2019qubit, ; li2020qubit, ; bravyi2022hybrid, ; bochniak2023quantum, ; salehi2022unconstrained, ; Ge_2020, ; cocchi2021graph, ; pichler2018quantum, ; kim2022rydberg, ; doern2007quantum, ; 8477865, ; pelofske2019solving, ; kerger2023asymptotically, ; pelofske2019solving, ; ramanujan2021approximate, ; kimmel2021query, ; witter2022query, ; dorn2009quantum, ; gabow1990data, ; mcleod2022benchmarking, ; vert2020revisiting, ; cameron2006quantum, ; bravyi2022hybrid, ; bravyi2020obstacles, ) to chemistry simulations (kassal2011simulating, ).
@@
Researchers are exploring how quantum computers can solve complex problems. They convert 3-SAT problems into a form that quantum computers can handle better. An improved method called ab-QAOA (yu2023solution, ) helps solve hard 3-SAT and Max-3-SAT problems using fewer steps. Quantum annealers, like D-Wave, are tested to solve SAT and MaxSAT problems. A bio-molecular (chang2021quantum, ) algorithm is proposed to solve the independent set problem efficiently. QAOA’s behaviour on graphs is studied, revealing its performance patterns. A quantum algorithm is developed to speed up solving the Traveling Salesman Problem (TSP). The Grover algorithm is used to tackle the Hamiltonian Cycle Problem (HCP) (jiang2022quantum, ).
@@
Researchers have been exploring quantum computing techniques to solve challenging optimization problems, such as the Multidimensional Knapsack Problem. They combined Quantum Particle Swarm Optimization with local search methods, yielding competitive outcomes. Another approach used the Quantum Approximate Optimization Algorithm (QAOA) to enhance solutions for constrained optimization problems, outperforming classical methods. The integration of the k-nearest neighbour technique and quantum cuckoo search led to better results in resource allocation and knapsack problems. A Diversity-Preserving Quantum Particle Swarm Optimization strategy (lai2020diversity, ) managed populations efficiently for improved solutions.
@@
The Quantum Alternating Operator Ansatz (cook2020quantum, ) offered approximate solutions for combinatorial problems, while the Quantum Approximate Optimization Algorithm tackled minimum vertex cover problems (pelofske2019solving, ; zhang2022applying, ; wang2023quantum, ; marsh2019quantum, ; wang2022classically, ). The Grover search algorithm achieved quantum speedup for the vertex cover problem. A Quantum-Walk assisted algorithm targeted bounded NP optimization problems. Notably, quantum algorithms were designed for matching problems, proving their effectiveness in various scenarios (jeffery2022multidimensional, ; bonnetain2020improved, ; mahasinghe2019solving, ; kempe2003quantum, ; childs2009universal, ; lovett2010universal, ; apers2019unified, ; marsh2019quantum, ; belovs2012learning, ).
Researchers have been exploring the capabilities of quantum computing in solving complex optimization problems. McLeod and Sasdelli (mcleod2022benchmarking, ) benchmarked D-Wave’s quantum annealers, Advantage, and 2000Q, using the maximum cardinality matching problem. They found that Advantage outperformed 2000Q for larger problems and analyzed the scaling of diabatic transition probability. Fowler addressed the minimum weight perfect matching problem on a modular qubit array, proving efficient solutions even in the presence of errors.
@@
In the realm of quantum walks, Jeffery and Zur (jeffery2022multidimensional, ) improved upper bounds on the time complexity of $k$-distinctness problems, offering insights into the quantum speedup, achieved in these scenarios.
@@
Harris and Kashefi (harris2023scalable, ) introduced an error mitigation protocol. The protocol is for running BQP-type computations on a quantum computer s.t. the noises are time-dependent. The authors hired verification techniques by Leichtle et al. (leichtle2021verifying, ). This allowed their results to interleave standard and test computation rounds. One should notice that the above technique provides a local correctness guarantee. In particular, it bounds the probability of returning a correct classical output exponentially. The authors further introduced a post-selection technique they named basketing. It could address time-dependent noise behaviours and also allow overhead reduction.
@@
3. Statistical Analysis
In this section, we provide a descriptive analysis of the quantum versus classical complexity of well-known computer science problems. We also demonstrate where different quantum approaches and solvers have been applied so far and could contribute to exploring future research directions. Later in the section, we draw remarks as an inferential analysis to help the researchers with their further investigations.
3.1. Descriptive Analysis
In this subsection, we provide a comparison, based on the existing results, between the complexity of quantum and classical approaches for well-known problems in computer science. We also demonstrate how various quantum approaches and solvers have been used for several problems, and refer to their potential future applications.
@@ 3.1. Descriptive Analysis
Problem Classic C. Quantum C. Improvement Problem Number(s) k-distinctness $\Theta(n\log n)$ $\tilde{\mathcal{O}}(n^{\frac{3}{4}-\frac{1}{4}\frac{1}{2^{k}-1}})$ Polynomial - (Element Distinctness) (In decision tree model (10.1145/800061.808735, ))  (jeffery2022multidimensional, ) k-SAT ($k\geq 3$) $\mathcal{O}(2^{1-\frac{\Theta(1)}{k}})$ (10.1145/3313276.3316359, ) $\mathcal{O}((2(1-1/k))^{n})$ (schoning1999probabilistic, ) - 1, 2, 15, 16, 17, 65 31, 32, 33, 34, 35 Maximum Independent Set $\mathcal{O}(1.1996^{n}\cdot n^{\mathcal{O}(1)})$ (XIAO2017126, ) $\mathcal{O}(1.1488^{n})$ (etde_21241168, ) Polynomial 66, 97, 103 Hamiltonian Cycle $\mathcal{O}^{*}(1.657^{n})$ (bjorklund2014determinant, ) $\mathcal{O}(n^{2n/(n+1)})$ (dorn2007quantum, ) Exponential 18, 10, 11, 14 Eulerian Tour $\mathcal{O}(|E|)$ (fleischner1990eulerian, ) $\mathcal{O}(\sqrt{n})$ (dorn2007quantum, ) Polynomial - Travelling Salesman Problem $\mathcal{O}(n^{2}\cdot 2^{n})$ (doi:10.1137/0110015, ) $\mathcal{O}({1.728}^{n}\cdot\text{poly}(n))$ (doi:10.1137/1.9781611975482.107, ) - 14 Maximum Matching $\mathcal{O}(|E|\sqrt{|V|})$ (micali1980v, ) $n^{3/2}\log^{2}n$ (beigi2022time, ) Polynomial 67, 85, 68, 70, 102 Maximum Cut $\text{exp}(n)$ (garey1997computers, ) $\sqrt{2^{n}/r}$ (chang2023quantum, ) - 8, 6 s.t. $r$ is num of max-cuts Graph Isomorphism $2^{\mathcal{O}(\log n)^{c}}$ (grohe2020graph, ) NA (alagic2007quantum, ) NA 95, 78, 30, 57 Maximum Clique $2^{n/4}$ (robson2001finding, ) $\mathcal{O}(\log_{2}n\cdot\sqrt{2^{n}})$ (haverly2021comparison, ) - 87, 88 String Edit Distance $\mathcal{O}(s^{2}+n)$ (landau1998incremental, ) $\mathcal{O}(n^{2}/\log n)$ (equi2021quantum, ) Poly-logarithmic 73, 90, 91 s.t. $s$ is a max. edit distance Longest Common Subsequence $\mathcal{O}(n\log^{2}n)$ (bhowmick2019approach, ) $\tilde{\mathcal{O}}(\sqrt{n})$ (le2023quantum, ) Polynomial 41 Protein Folding $\omega(n^{c}),\ \forall c>0$ $\mathcal{O}(2^{n/2})$ (wong2022fast, ) - 27, 100, 101 Mixed-Integer Programming $(\log n)^{{}^{\mathcal{O}(n)}}\cdot\text{poly}(n)$ (reis2023subspace, ) $\exp(n)$ 13 Knapsack Problem $\mathcal{O}(2^{n/2})$ (doi:10.1137/0210033, ) $\mathcal{O}(2^{n/2})$ (fujimura2010quantum, ) - 38, 37, 6.12, 40 Subset-Sum $\tilde{\mathcal{O}}(2^{0.291n})$ (becker2011improved, ) $\tilde{\mathcal{O}}(2^{0.236n})$ (bonnetain2020improved, ) - 36, 43, 44, 45 Discrete Logarithm $\omega(n^{c}),\ \forall c>0$ $\mathcal{O}(\log|\mathcal{G}|)$ (hhan2023quantum, ) Superpolynomial - for a cyclic group $\mathcal{G}$ of prime order Integer Factorization $\exp(1.9(\log n)^{\frac{1}{3}}(\log\log n)^{\frac{2}{3}})$ (buhler1993factoring, ) ${\mathcal{O}}(b^{3})$ (shor1999polynomial, ) Superpolynomial - $n$ is the number to factor s.t. $b$ is num of bits of $n$ Crypto. Certified Random Bits - $\exp(n)$ (aaronson2023certified, ) - 49, 54